OpenStudy (anonymous):

Find the equation of tangent and normal line to the given curves.. 1.) x^2 + y^2 = 25; (-3,-4) 2.) x^2/32 + y^2/2 =1; (-4,1) 3.) x^2 -6x+2y-8=0; x=0

OpenStudy (anonymous):

Same step as the other topic. You need to solve it in order to "y" then you need to derivate the given curves then you need to use this formula: y=f′(a)(x−a)+f(a) and now you have the equation of the tangent. Something like y=mx+b To find the normal line you just need to do m=1/m

OpenStudy (anonymous):

hmmm maybe not

OpenStudy (anonymous):

for 1 take the derivative and get \[2x+2yy'=0\] \[y'=-\frac{x}{y}\]

OpenStudy (anonymous):

so at \[(-3,-4)\] the slope is \[-\frac{3}{4}\] and equation of tangent line is \[y+4=-\frac{3}{4}(x+3)\] equation of normal line is \[y+4=\frac{4}{3}(x+3)\]

OpenStudy (anonymous):

second one i cannot read

OpenStudy (anonymous):

it means Xsquared over 32 + ysquared over 2 =1

OpenStudy (anonymous):

Ellipse is same as circle, implicit diff/normal etc and last is just a normal quadratic.